find the nth term of an A.P. whose (r+1)th term is 4r-3
Answer by stanbon(48568)

find the nth term of an A.P. whose (r+1)th term is 4r-3
------------------------------
Formula: a(n) = a(1) + (n-1)d
----------------------------------
Your Problem:
If r+1 = 1, r = 0
Therefore a(1) = 4*0-3 = -3
-----------------------------------
If r+1 = 2, r = 1
Therefore a(2) = 4*1-3 = 1
-----------------------------------
So d = a(2)-a(1) = 1-(13) = 4
=-----------------------------------
Your Problem:
To find the nth term:
If r+1 = n, r = n-1
Therefore a(n) = 4(n-1)-3 = 4n-7
=========================================
Cheers,
Stan H.